# LAG: Logic-Augmented Generation from a Cartesian Perspective

### v2

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, yet exhibit critical limitations in knowledge-intensive tasks, often generating hallucinations when faced with questions requiring specialized expertise. While retrieval-augmented generation (RAG) mitigates this by integrating external knowledge, it struggles with complex reasoning scenarios due to its reliance on direct semantic retrieval and lack of structured logical organization. Inspired by Cartesian principles from Discours de la méthode, this paper introduces Logic-Augmented Generation (LAG), a novel paradigm that reframes knowledge augmentation through systematic question decomposition and dependency-aware reasoning. Specifically, LAG first decomposes complex questions into atomic sub-questions ordered by logical dependencies. It then resolves these sequentially, using prior answers to guide context retrieval for subsequent sub-questions, ensuring stepwise grounding in logical chain. To prevent error propagation, LAG incorporates a logical termination mechanism that halts inference upon encountering unanswerable sub-questions and reduces wasted computation on excessive reasoning. Finally, it synthesizes all sub-resolutions to generate verified responses. Experiments on four benchmark datasets demonstrate that LAG significantly enhances reasoning robustness, reduces hallucination, and aligns LLM problem-solving with human cognition, offering a principled alternative to existing RAG systems.
Introduction
@@
Large language models (LLMs), like Claude (Anthropic 2024), ChatGPT (OpenAI 2023) and the Deepseek series (Liu et al. 2024), have demonstrated remarkable capabilities in many real-world tasks (Chen et al. 2024b; Zhou et al. 2025), such as question answering (Allam and Haggag 2012), text comprehension (Wright and Cervetti 2017) and content generation (Kumar 2024). Despite the success, these models are often criticized for their tendency to produce hallucinations, generating incorrect statements on tasks beyond their knowledge and perception (Ji et al. 2023; Zhang et al. 2024). Recently, retrieval-augmented generation (RAG) (Gao et al. 2023; Lewis et al. 2020) has emerged as a promising solution to alleviate such hallucinations. By dynamically leveraging external knowledge from textual corpora, RAG enables LLMs to generate more accurate and reliable responses without costly retraining (Lewis et al. 2020; Devalal and Karthikeyan 2018).
Figure 1: Comparison of three paradigms. LAG exhibits greater lightweight properties compared to GraphRAG while achieving higher accuracy than RAG and GraphRAG.
@@
RAG systems typically operate through three key stages: knowledge preprocessing, retrieval, and integration. First, external textual corpora are segmented into manageable chunks and converted into vector representations for efficient indexing. When a query is received, the system then retrieves relevant text segments using semantic similarity matching (Sawarkar, Mangal, and Solanki 2024) or keyword-based search (Purwar and Sundar 2023). Finally, during integration, the retrieved information is combined with the original query to produce knowledge-enhanced responses. Recent advances in RAG technology have evolved beyond basic text retrieval toward more sophisticated approaches. These include graph-based systems (Zhang et al. 2025; Peng et al. 2024; Procko and Ochoa 2024; Shengyuan et al. 2023) that model conceptual relationships using graph structures, hierarchical methods (Chen et al. 2024a; Li et al. 2025b; Chen et al. 2025) preserving document organization through multi-level retrieval, re-ranking implementations (Glass et al. 2022; Xu, Shi, and Choi 2024) utilizing preliminary retrieval followed by refined scoring, Self-RAG architectures (Asai et al. 2024) capable of on-demand retrieval and self-reflection, and adaptive frameworks (Tang et al. 2025; Sarthi et al. 2024) that dynamically adjust retrieval strategies based on query complexity. These strategies significantly enhance naive RAG systems through improved retrieval accuracy.
@@
To bridge this gap, we introduce Logic-Augmented Generation (LAG), a novel paradigm inspired by Cartesian principles outlined in Discours de la méthode. LAG introduces a new reasoning-first pipeline that integrates systematic decomposition and controllable reasoning into retrieval-augmented generation. Instead of immediately invoking a retriever, LAG begins by carefully analyzing the question and breaking it down into a set of atomic sub-questions that follow a logical dependency structure. The system then answers these sub-questions step by step, starting with the most basic ones that do not rely on any prior information. As each sub-question is resolved, its answer becomes part of the context used to guide the retrieval and resolution of the next, more dependent sub-question. The final answer is synthesized only after all necessary sub-questions have been addressed. If an inconsistency arises during reasoning, the logical terminator is triggered to activate the alternative solution. Generally, our main contribution is listed as follows:
  • We identify the limitations of existing RAG systems when dealing with complex questions, and propose LAG, a new reasoning-first pipeline that integrates systematic decomposition and logical reasoning into RAG systems.
  • LAG first decomposes complex questions into atomic sub-questions ordered by logical dependencies. It then resolves these sequentially, using prior answers to guide context retrieval for subsequent sub-questions, ensuring stepwise grounding in logical chain.
@@
Related Work
@@
RAG has emerged as a critical framework for enhancing LLMs by integrating external knowledge. Early approaches such as REALM (Guu et al. 2020) and DPR (Karpukhin et al. 2020) focus on encoding large text corpora into dense embeddings. In recent years, GraphRAG has become a new direction because it can structure fragmented knowledge. RAPTOR (Sarthi et al. 2024) and Microsoft’s GraphRAG (Edge et al. 2025) both use hierarchical clustering: RAPTOR constructs recursive trees with multi-level summarization, and GraphRAG applies community detection with LLM-generated synopses, to support coarse-to-fine retrieval and high-coverage responses. DALK (Li et al. 2024) and KGP (Wang et al. 2024) introduce dynamic KG construction and traversal agents, using LLMs to build domain-specific graphs and self-aware retrieval policies, to inject structural context while reducing noise. GFM-RAG (Luo et al. 2025), G-Retriever (He et al. 2024), and LightRAG (Guo et al. 2025) combine graph neural encoders with specialized retrieval objectives: a query-dependent GNN trained in two stages for multi-hop generalizability, a Prize Collecting Steiner Tree formulation to reduce hallucination and improve scalability, and a dual-level graph-augmented index for efficient, incrementally updatable lookup, respectively enabling accurate, scalable reasoning over graphs. HippoRAG (Gutiérrez et al. 2024), inspired by hippocampal memory processes, leverages Personalized PageRank for single-step multi-hop retrieval, delivering state-of-the-art efficiency and performance on both path following and path finding QA tasks. HippoRAG2 (Gutiérrez et al. 2025) further optimizes knowledge graph refinement and deeper passage integration. More details regarding related work are provided in the Appendix A.
Figure 2: The overall framework of our proposed LAG.
Preliminaries.
@@
Retrieval-Augmented Generation (RAG) enhances language models by incorporating external knowledge retrieved from a large corpus. We denote the input as a natural language question $q$, which may involve latent constraints, or multi-hop reasoning. The system has access to a retrieval corpus $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$, where each $c_{i}$ represents a passage, document chunk, or knowledge entry consisting of unstructured text. These entries may vary in granularity and source (e.g., Wikipedia, scientific papers, web documents), but are assumed to be independently indexable and retrievable. Given a query $q$ or any intermediate sub-question $q^{\prime}$, a retriever $\mathcal{R}$ returns a ranked list of relevant passages $\mathcal{R}(q^{\prime})\subset\mathcal{C}$ to support downstream reasoning. Each retrieved item $c\in\mathcal{C}$ is treated as a semantically self-contained unit of information, which the system uses as external evidence during the generation or verification process.
The Framework of LAG
@@
In Discours de la méthode, Descartes proposed four principles for solving problems scientifically: 1) Doubt everything, avoiding precipitancy and prejudice. 2) Divide any complex question into multiple simpler subquestions. 3) Order sub‑questions from the simplest to the most complex and fix them step by step. 4) Once all issues are solved, review them to ensure that nothing was omitted. Inspired by this principle, LAG introduces a novel reasoning‑first paradigm that directly aligns with it. First, to avoid precipitancy, LAG does not perform direct retrieval of the entire question. Second, the adaptive decomposition module decomposes the complex query into multiple atomic sub-questions. Third, the logical reorder module arranges these sub-questions according to their logical dependency, and the logical chain reasoning module resolves them accordingly. Finally, final answers are constructed by logically combining all sub-solutions, followed by validation against the original question to ensure complete coverage. Notably, a insurance mechanism is incorporated into this process: when the reasoning process is deemed invalid, the logical terminator is activated to initiate the alternative solution.
Adaptive Question Decomposition
@@
where the Cognitive Load metric $\mathrm{CL}(q)$ integrates:
$\mathrm{CL}(q)=\underbrace{\sigma(\mathrm{Var}(\phi(q)))}_{\textsc{Semantic Scope}}+\underbrace{\sigma(\mathrm{Depth}(q))}_{\textsc{Reasoning Steps}}+\underbrace{\sigma(\mathcal{H}(q))}_{\textsc{Ambiguity}}$ (2)
The Cognitive Load metric $\mathrm{CL}(q)$ integrates three complementary signals to estimate the complexity of a question. The first term, Semantic Scope, is computed as the variance of the question’s embedding $\phi(q)$, capturing how broad the question is semantically. A higher value often indicates a wider topic range or more entangled concepts. The second term, Reasoning Steps, measures the depth of compositional reasoning required to answer $q$. We estimate this by counting the number of latent inference steps involved. The final term, Ambiguity, quantifies semantic uncertainty through a heuristic entropy-based function $\mathcal{H}(q)$, which reflects referential ambiguity (e.g., unclear pronouns). $\sigma(\cdot)$ represents normalization function. Once $\mathrm{CL}(q)$ exceeds the threshold $\tau(t)$, which decays over time to encourage early resolution, our module recursively fractures $q$ into smaller sub-questions until all resulting $q_{i}$ satisfy $\mathrm{CL}(q_{i})\leq\tau(t)$. This recursive refinement balances the need for logical soundness and factual verifiability with the goal of minimizing unnecessary conversational turns.
Logical Chain Reasoning
@@
Logic-Guided Retrieval
After prior sub-question is answered, we update the retrieval query by incorporating both its corresponding answer and the subsequent sub-question into a single textual context. Instead of directly combining embeddings, we concatenate the prior answer $a_{i}$ and subsequent sub-question $q_{i+1}$ into a natural language form, e.g., $A_{i}:a_{i}$, $Q_{i+1}:q_{i+1}$”, and encode the resulting text to obtain a query vector for the next retrieval step. Formally, the query embedding at step $i+1$ is computed as:
@@
where $\texttt{concat}(q_{i},a_{i})$ denotes the textual concatenation of the sub-question and prior answer, and $\phi(\cdot)$ is a shared encoder used for both questions and passages. The vector $\mathbf{q}^{(i+1)}$ is then used to query the corpus $\mathcal{C}$, retrieving a set of passages $\mathcal{R}(q^{(i+1)})$ to support resolution of the sub-question $q_{i+1}$. This context-aware retrieval process allows the system to progressively incorporate verified knowledge into subsequent steps, enabling more precise and contextually grounded evidence collection along the logical chain.
Logical Terminator
To ensure both efficiency and robustness, we design an automatic stopping mechanism that prevents excessive or unnecessary reasoning during the logical chain reasoning. This component plays a critical role in avoiding error propagation from unanswerable sub-questions and reducing wasted computation on low-value expansions. By monitoring retrieval confidence, logical dependency states, and semantic redundancy, the system dynamically determines when to halt further reasoning, ensuring that the model focuses its efforts only when informative progress can be made.
❶ Retriever Confidence Drop. We monitor the retriever’s output quality across consecutive sub-questions. Specifically, if the top-$k$ retrieved passages for a given sub-question all exhibit low semantic similarity with the query embedding, the system interprets this as a signal of insufficient external support. Let $\text{sim}(\mathbf{q}^{\prime},c_{i})$ denote the cosine similarity between a sub-question $q^{\prime}$ and a retrieved passage $c_{i}\in\mathcal{R}(q^{\prime})$. If
@@
❷ Dependency Exhaustion. In our logical chain reasoning framework, sub-questions are arranged based on their logical dependencies. When all prerequisite sub-questions for query $q^{\prime}$ have been successfully resolved, but the query still lacks sufficient support or a valid answer, the system considers the reasoning chain to be exhausted. Formally, let $\text{Deps}(q^{\prime})=\{q_{1},q_{2},\dots,q_{m}\}$ denote the set of sub-questions that $q^{\prime}$ depends on. If all $q_{i}\in\text{Deps}(q^{\prime})$ are answered and yet $q^{\prime}$ cannot be resolved, we halt further reasoning:
@@
❸ Semantic Saturation and Step Limit. To prevent excessive reasoning depth, we introduce two additional halting conditions. First, semantic saturation is detected when the newly retrieved passages show high redundancy with the accumulated context. Let $\mathcal{C}_{\text{prev}}$ denote previously retrieved content, and $c_{\text{new}}$ a candidate passage.
$\text{sim}(c_{\text{new}},\mathcal{C}_{\text{prev}})>\gamma,$ (6)
For most $c_{\text{new}}$, where $\gamma$ is a redundancy threshold (0.9), the system considers the information space saturated. Second, to guarantee tractability, we define a maximum number of reasoning steps $T_{\max}$ (5). Once the number of resolved sub-questions reaches $T_{\max}$, the reasoning process is terminated.
Integrated Generation
As shown in Figure 2, our framework synthesizes the final answer by integrating all validated sub-question responses through a composition process. The system retrieves relevant evidence for each sub-query. Building upon the established reasoning chain, we first generate a comprehensive draft answer that incorporates each verified sub-solution while maintaining logical coherence with the original query. This draft must properly address all sub-questions without contradiction while fully covering the scope of the initial problem. When inconsistencies are identified, the logical terminator will stop further reasoning and merely retain the reliable logical chain, then proceed to the alternative solution, which will feed sub-questions, reliable logical chain, and retrieved context to the LLM to generate the final response. For unresolved sub-questions, instead of relying on prior knowledge, the information retrieved will be used as the basis. This mechanism maintains the natural flow of human-like explanation rather than simply concatenating facts.

### v3

@@ Abstract
Large language models (LLMs) have demonstrated remarkable capabilities across a wide range of tasks, yet exhibit critical limitations in knowledge-intensive tasks, often generating hallucinations when faced with questions requiring specialized expertise. While retrieval-augmented generation (RAG) mitigates this by integrating external knowledge, it struggles with complex reasoning scenarios due to its reliance on direct semantic retrieval and lack of structured logical organization. Inspired by Cartesian principles from Discours de la méthode, this paper introduces Logic-Augmented Generation (LAG), a novel paradigm that reframes knowledge augmentation through systematic question decomposition, atomic memory bank and logic-aware reasoning. Specifically, LAG first decomposes complex questions into atomic sub-questions ordered by logical dependencies. It then resolves these sequentially, using prior answers to guide context retrieval for subsequent sub-questions, ensuring stepwise grounding in the logical chain. Experiments on four benchmarks demonstrate that LAG significantly improves accuracy and reduces hallucination over existing methods.
LAG: Logic-Augmented Generation from a Cartesian Perspective
Yilin Xiao, Chuang Zhou, Yujing Zhang, Qinggang Zhang Su Dong, Shengyuan Chen, Chang Yang, Xiao Huang The Hong Kong Polytechnic University Hong Kong SAR yilin.xiao@connect.polyu.hk, {qinggang.zhang, xiao.huang}@polyu.edu.hk
1 Introduction
@@
Large language models (LLMs), like Claude Anthropic (2024), ChatGPT OpenAI (2023) and the Deepseek series Liu et al. (2024), have demonstrated remarkable capabilities in many real-world tasks, such as question answering Allam and Haggag (2012), text comprehension Wright and Cervetti (2017) and content generation Kumar (2024). Despite the success, these models are often criticized for their tendency to produce hallucinations, generating incorrect statements on tasks beyond their perception Ji et al. (2023); Zhang et al. (2024). Recently, retrieval-augmented generation (RAG) Gao et al. (2023); Lewis et al. (2020) has emerged as a promising solution to alleviate such hallucinations. By dynamically leveraging external knowledge from textual corpora, RAG enables LLMs to generate more accurate and reliable responses without costly retraining Lewis et al. (2020); Devalal and Karthikeyan (2018).
Figure 1: Comparison of three RAG paradigms. LAG offers a superior balance of efficiency and accuracy, providing a more lightweight solution than GraphRAG while outperforming it and traditional RAG in accuracy.
@@
RAG systems typically operate through three key stages: knowledge preprocessing, retrieval, and integration. First, external textual corpora are segmented into manageable chunks and converted into vector representations for efficient indexing. When a query is received, the system then retrieves relevant text segments using semantic similarity matching Sawarkar et al. (2024) or keyword-based search Purwar and Sundar (2023). Finally, during integration, the retrieved information is combined with the original query to produce knowledge-enhanced responses. Recent advances in RAG technology have evolved beyond basic text retrieval toward more sophisticated approaches. These include graph-based systems Zhang et al. (2025); Peng et al. (2024); Procko and Ochoa (2024) that model conceptual relationships using graph structures, hierarchical methods Chen et al. (2024); Li et al. (2025b) preserving document organization through multi-level retrieval, re-ranking implementations Glass et al. (2022); Xu et al. (2024) utilizing preliminary retrieval followed by refined scoring, Self-RAG architectures Asai et al. (2024) capable of on-demand retrieval and self-reflection, and adaptive frameworks Tang et al. (2025); Sarthi et al. (2024) that dynamically adjust retrieval strategies based on query complexity. These strategies significantly enhance naive RAG systems through improved retrieval accuracy.
@@
To bridge this gap, we introduce Logic-Augmented Generation (LAG), a novel paradigm inspired by Cartesian principles outlined in Discours de la méthode. LAG introduces a new reasoning-first pipeline that integrates systematic decomposition, atomic memory bank and controllable reasoning into retrieval-augmented generation. Instead of immediately invoking a retriever, LAG begins by carefully analyzing the question and breaking it down into a set of atomic sub-questions that follow a logical dependency structure. To enhance efficiency, LAG utilizes an atomic memory bank, which stores and retrieves cached, high-confidence solutions to recurrent atomic sub-questions. The system then answers these sub-questions step by step, first checking the memory bank for verified solutions to avoid redundancy and potential hallucinations. When a novel sub-question is encountered, it is resolved from first principles. This process starts with the most basic, independent sub-questions. As each sub-question is resolved, its answer becomes part of the context used to guide the retrieval and resolution of the next, more dependent sub-question. The final answer is synthesized only after all necessary sub-questions have been addressed. If an inconsistency arises during reasoning, the logical terminator triggers the activation of the alternative solution. Our main contribution is listed as follows:
  • We identify key limitations of RAG, and propose LAG, a reasoning-first framework that integrates systematic decomposition, atomic memory bank and logical reasoning.
  • LAG decomposes questions into logically-dependent sub-questions and resolves them sequentially, first retrieving verified solutions from the memory bank, then solving novel sub-questions following the logical structure.
@@
Figure 2: The framework of LAG. (I) Adaptive question decomposition splits complex queries into atomic sub-questions using a cognitive load. (II) Logical chain reasoning resolves sub-Q by the order of logical dependency, utilizing an atomic memory bank for recurrent knowledge. Logical terminator halts unreliable chains early. Finally, answers are synthesized via integrated generation (complete chains) or alternative solution (terminated chains).
2 Related Work
@@
RAG has emerged as a critical framework for enhancing LLMs by integrating external knowledge. Early approaches such as REALM Guu et al. (2020) and DPR Karpukhin et al. (2020) focus on encoding large text corpora into dense embeddings. In recent years, GraphRAG has become a new direction because it can structure fragmented knowledge. RAPTOR Sarthi et al. (2024) and Microsoft’s GraphRAG Edge et al. (2025) both use hierarchical clustering: RAPTOR constructs recursive trees with multi-level summarization, and GraphRAG applies community detection with LLM-generated synopses, to support coarse-to-fine retrieval and high-coverage responses. DALK Li et al. (2024) and KGP Wang et al. (2024) introduce dynamic KG construction and traversal agents, using LLMs to build domain-specific graphs and self-aware retrieval policies, to inject structural context while reducing noise. GFM-RAG Luo et al. (2025), G-Retriever He et al. (2024), and LightRAG Guo et al. (2025) combine graph neural encoders with specialized retrieval objectives: a query-dependent GNN trained in two stages for multi-hop generalizability, a Prize Collecting Steiner Tree formulation to reduce hallucination and improve scalability, and a dual-level graph-augmented index for efficient, incrementally updatable lookup, respectively enabling accurate, scalable reasoning over graphs. HippoRAG Gutiérrez et al. (2024), inspired by hippocampal memory processes, leverages Personalized PageRank for single-step multi-hop retrieval, delivering state-of-the-art efficiency and performance on both path following and path finding QA tasks. HippoRAG2 Gutiérrez et al. (2025) further optimizes knowledge graph refinement and deeper passage integration. More details of related work are provided in the Appendix.
3 Preliminaries.
@@
Retrieval-Augmented Generation (RAG) enhances language models by incorporating external knowledge retrieved from a large corpus. We denote the input as a natural language question $q$, which may involve latent constraints, or multi-hop reasoning. The system has access to a retrieval corpus $\mathcal{C}=\{c_{1},c_{2},\dots,c_{N}\}$, where each $c_{i}$ represents a passage, document chunk, or knowledge entry consisting of unstructured text. These entries may vary in granularity and source (e.g., Wikipedia, scientific papers, web documents), but are assumed to be independently indexable and retrievable. Given a query $q$, a retriever $\mathcal{R}$ returns a ranked list of relevant passages $\mathcal{R}(q)\subset\mathcal{C}$ to support downstream reasoning. Each retrieved item $c\in\mathcal{C}$ is treated as a semantically self-contained unit of information, which the system uses as external evidence during the generation or verification process.
4 The Framework of LAG
@@
In Discours de la méthode, Descartes proposed four principles for solving problems scientifically: (i) Doubt everything, avoiding precipitancy and prejudice. (ii) Divide any complex question into multiple simpler subquestions. (iii) Order sub‑questions from the simplest to the most complex and fix them step by step. (iv) Once all issues are solved, review them to ensure that nothing was omitted. Inspired by this principle, LAG introduces a novel reasoning‑first paradigm that directly aligns with it. First, to avoid precipitancy, LAG does not perform direct retrieval of the entire question. Second, the adaptive decomposition module decomposes the complex query into multiple atomic sub-questions. Third, the logical reorder module arranges these sub-questions according to their logical dependency, and the logical chain reasoning module resolves them accordingly. Finally, final answers are constructed by logically combining all sub-solutions, followed by validation against the original question to ensure complete coverage. Notably, to enhance both efficiency and robustness, LAG incorporates two key mechanisms into its reasoning pipeline: (i) a logic-guided reasoning strategy that leverages resolved sub-questions to guide subsequent retrieval and answering, and (ii) an insurance mechanism where a logical terminator is activated to initiate the alternative solution pathway if the reasoning process is deemed invalid.
4.1 Adaptive Question Decomposition
@@
where the Cognitive Load metric comprises:
$\mathrm{CL}(q)=\underbrace{\sigma(\mathrm{Var}(\phi(q)))}_{\textsc{Semantic Scope}}+\underbrace{\sigma(\mathrm{Depth}(q))}_{\textsc{Reasoning Steps}}+\underbrace{\sigma(\mathcal{H}(q))}_{\textsc{Ambiguity}}$
The Cognitive Load metric $\mathrm{CL}(q)$ integrates three complementary signals to estimate the complexity of a question. Firstly, Semantic Scope, is computed as the variance of the question’s embedding $\phi(q)$, capturing how broad the question’s semantic coverage is. A higher value often indicates a wider topic range or more entangled concepts. The second term, Reasoning Steps, measures the depth of compositional reasoning required to answer $q$. LLMs estimate this by counting the number of latent inference steps involved. Lastly, Ambiguity, quantifies semantic uncertainty through a heuristic entropy-based function $\mathcal{H}(q)$, which reflects referential ambiguity (Details are in the appendix). $\sigma(\cdot)$ represents normalization function. Once $\mathrm{CL}(q)$ exceeds the threshold $\tau(t)$, which decays over time to encourage early resolution, our module recursively fractures $q$ into smaller sub-questions until all resulting $q_{i}$ satisfy $\mathrm{CL}(q_{i})\leq\tau(t)$. This recursive refinement balances the need for logical soundness and factual verifiability with the goal of minimizing unnecessary conversational turns.
4.2 Logical Chain Reasoning
@@
4.2.1 Logic-Guided Retrieval
After prior sub-question is answered, we update the retrieval query by incorporating both its corresponding answer and the subsequent sub-question into a single textual context. Instead of directly combining embeddings, we concatenate the prior answer $a_{i}$ and subsequent sub-question $q_{i+1}$ into a natural language form, e.g., $A_{i}:a_{i}$, $Q_{i+1}:q_{i+1}$”, and encode the resulting text to obtain a query vector for the next retrieval step. Formally, the query embedding at step $i+1$ is shown as:
@@ 4.2.1 Logic-Guided Retrieval
where $\texttt{concat}(q_{i},a_{i})$ denotes the textual concatenation of the sub-question and prior answer, and $\phi(\cdot)$ is a shared encoder used for both questions and passages. The vector $\mathbf{q}^{(i+1)}$ is then used to query the corpus $\mathcal{C}$, retrieving a set of passages $\mathcal{R}(q^{(i+1)})$ to support resolution of the sub-question $q_{i+1}$. This context-aware retrieval process allows the system to progressively incorporate verified knowledge into subsequent steps, enabling more precise evidence collection along the logical chain.
4.2.2 Logical Terminator
To ensure both efficiency and robustness, we design an automatic stopping mechanism that prevents excessive or unnecessary reasoning during the logical chain reasoning. This component plays a critical role in avoiding error propagation from unanswerable sub-questions and unnecessary computation. By monitoring retrieval confidence, logical dependency states, and semantic redundancy, the system dynamically determines when to halt further reasoning, ensuring that the model focuses its efforts only when informative progress can be made.
Retriever Confidence Drop. We monitor the retriever’s output across consecutive sub-questions. If the top-$k$ retrieved passages for a given sub-question all exhibit low semantic similarity with the query, the system interprets this as a signal of insufficient external support. Let $\text{sim}(\mathbf{q}^{\prime},c_{i})$ denote the cosine similarity between a sub-question $q^{\prime}$ and a retrieved passage $c_{i}\in\mathcal{R}(q^{\prime})$. If
@@ 4.2.2 Logical Terminator
Dependency Exhaustion. In our logical chain reasoning framework, sub-questions are arranged based on their logical dependencies. When all prerequisite sub-questions for query $q^{\prime}$ have been successfully resolved, but the query still lacks sufficient support or a valid answer, the system considers the reasoning chain to be exhausted. Formally, let $\text{Deps}(q^{\prime})=\{q_{1},q_{2},\dots,q_{m}\}$ denote the set of sub-questions that $q^{\prime}$ depends on. If all $q_{i}\in\text{Deps}(q^{\prime})$ are answered and yet $q^{\prime}$ cannot be resolved, we halt further reasoning:
@@
Semantic Saturation. Beyond static step limits defined in the previous section, our framework actively monitors reasoning progress and halts retrieval based on a semantic saturation criterion. This is determined by evaluating the redundancy of new information relative to the accumulated context. Let $\mathcal{C}_{\text{prev}}$ denote the set of previously retrieved passages and $c_{\text{new}}$ a candidate passage. The framework deems the information space saturated and halts retrieval when the similarity $\text{sim}(c_{\text{new}},\mathcal{C}_{\text{prev}})>\gamma$ ( $\gamma$ is set to be 0.9). This prevents the system from iterating further when little new information is being uncovered.
4.2.3 Atomic Memory Bank
To enhance reasoning consistency and efficiency, we introduce the Atomic Memory Bank, designed to store atomic knowledge units. By caching and reusing high-confidence solutions to atomic sub-question (generated via LAG), we mitigate the redundancy and hallucinations associated with repeatedly invoking the LLM for identical sub-tasks.
Formally, the memory bank is maintained as $\mathcal{M}=\{(q^{\prime}_{i},a^{\prime}_{i},\mathbf{q^{\prime}}_{i})\}_{i=1}^{|\mathcal{M}|}$, comprising tuples of a sub-question $q^{\prime}_{i}$, its answer $a^{\prime}_{i}$, and the vector embedding $\mathbf{q^{\prime}}_{i}$. During the inference, for a given query sub-question $q^{\prime}_{*}$, we compute its embedding $\mathbf{q^{\prime}}_{*}$ and retrieve the stored answer if the similarity with an existing entry exceeds a threshold $\gamma$:
$\max_{i}\frac{\mathbf{q^{\prime}}_{*}\cdot\mathbf{q^{\prime}}_{i}}{\|\mathbf{q^{\prime}}_{*}\|\|\mathbf{q^{\prime}}_{i}\|}>\gamma,$ (5)
where the threshold $\gamma$ is set to 0.9. This value coincides with the semantic saturation threshold, establishing a unified criterion for the system’s critical decisions. For the memory update process, we implement a strict filtering protocol to ensure knowledge quality. When the LLM solves a new sub-question, it assigns a confidence score on a five-point scale. Only pairs achieving the highest confidence level are candidates for storage. Furthermore, we incorporate a Context Independence Validation to detect and filter sub-questions containing anaphoric references or other context-dependent ambiguities, which could be found at Appendix. Only sub-questions that pass both the confidence and independence checks are committed to $\mathcal{M}$.
4.3 Integrated Generation
As shown in Figure 2, our framework synthesizes the final answer by integrating all validated sub-question responses through a composition process. The system retrieves relevant evidence for each sub-query. Building upon the established reasoning chain, we first generate a comprehensive draft answer that incorporates each verified sub-solution while maintaining logical coherence with the original query. This draft must properly address all sub-questions without contradiction while fully covering the scope of the initial problem. When inconsistencies are identified, the logical terminator will stop further reasoning and merely retain the reliable logical chain, then proceed to the alternative solution, which will feed sub-questions, reliable logical chain, and retrieved context to the LLM to generate the final response. For unresolved sub-questions, instead of relying on prior knowledge, the information retrieved will be used as the basis.